HUBBLE SPACE TELESCOPE DAILY REPORT #907 PERIOD COVERED: 0800 EST Daily Status Report as of 165/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Four Sets of PC 4714 (WF/PC Cycle 2 Calibration: Internal Delta Flats) The Planetary Camera was used to make a large series of internal delta flat observations. These observations will be used for the restoration fo Earth flats, which were altered due to changes in quantum efficiency which occurred after WFPC decontamination operations. All of the observations were monitored in the OSS. They were reported to look okay. 1.2 Completed Three Sets of WFC 4848 (GTO Low Latitude Parallel Program Cycle 3 High) A grand total of sixteen low latitude sky survey images were made with the Wide Field Camera and were taken in parallel with the prime science observations. All sixteen of the images were reported to look normal for this type of parallel observations. One observation showed a moving object which trailed across its image. 1.3 Completed Two Sets of FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) The Faint Object Camera was used to complete one set of snapshot ultraviolet observations of nearby galaxy ESO-0052-3. The exposure lasted for ten minutes and used filter F220W. The single observations was received and displayed in the OSS. It was reported to look normal. The second scheduled observation with the Faint Object Camera did no occur because the spacecraft was in internal hold Safemode. (See Note 2.3) 1.4 Completed WFC 4929 (WF/PC Cycle 2 : Improved Dark Calibration) The Wide field Camera was used to make two Dark Calibration exposures. The exposures were 1200 seconds long, and the camera was pre-flashed. They both were reported to look completely normal. 1.5 Completed FOS/BL 4697 (FOS Polarimetric Calibration) The Faint Object Spectrograph was used on both sides (i.e., Red and Blue Sides) to obtain a series of measurements of the internal Pt-Ne lamps to locate the two oppositely polarized spectra that occur with the polarimeter. The observations were executed as scheduled, and no problems were reported. 1.6 Completed WFC 4277 (Calibration of Nearby Galaxy Type IA Supernovae as Standard Candles) The Wide Field Camera was used to make Two images of the galaxy NGC-5253 to determine the periods of Cepheid variables in this galaxy. This is hoped to make it possible to accurately determine the distance to this galaxy. Both observations were received and displayed. The looked okay, although the core region of one image was saturated. 1.7 Completed WFC/PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) Rapid Internal Monitoring observations were made with both the Wide Field Camera and the Planetary Camera using filter F555W to look for new forms of contamination (i.e. Daisies and Measles), and the possible return of quantum efficiency hysteresis. Both image were received and displayed. They both were reported to look okay. 1.8 Completed FOS/RD 4062 (Stellar and Gas Dynamics in Normal Galaxies Augmentation: Cycle 2) The Faint Object Spectrograph (Red Side) was used to make a series of spectral observations over the inner arcsecond of the central region of galaxy NGC-224 (aka, M-31 the Andromeda galaxy). The observations were made on a rectangular grid at positions having a spacing of 0.25 arcseconds. The wavelenght calibration observations was normal. The confirmation image showed the target well centered in the aperture. All of the spectral observations were reported to look normal. One of the spectral observations was reported to have had a recentering event occur during the observation. 1.9 Completed FOC/96 3801 (Search for QSOs Suitable for Subsequent Observation fo He II) The Faint Object Camera was used to make to make one observation of quasar 0053-303 to measure the levels of its ultraviolet flux, to see if the flux levels are high enough for subsequent observations using the GHRS and FOS. The observations was received and displayed in the OSS. No anomalies were noted. 1.10 Completed GHRS/2 4066 (GHRS Echelle Spectral Lamp Mini- Function Test) The High Resolution Spectrograph (Echelle Mode) was used to conduct a mini-lamp test for an internal test and calibration. The Spectral lamp SC2 was used with the Echelle B spectrograph to calibrate the carrousel function, Y deflections, resolving power, sensitivity, and scattered light. The OSS monitored the observations, and assessed the data quality. No problems were reported. 1.11 Completed GHRS/2 4065 (Dark Noise Monitoring HRS) The High Resolution spectrograph was used on a strictly internal test for routine monitoring of detector dark noise. The observations were executed as scheduled, and no problems were reported. 1.12 Completed PC 4764 (Investigation of the Dynamics of Binary Asteroids) The Planetary Camera was used to make four images of the Asteroid Hektor. This asteroid is suspected of being a physical contact binary asteroid because of its unusual light curves. All four observations were received and displayed. They were reported to look okay. 1.13 Completed FOS/BL 3441 (Post Asymptotic Giant Branch Evolution in the Magellanic Clouds) The Faint Object Spectrograph was used to make spectral observations of the planetary nebula SMC-SMP69 which is located in the Small Magellanic Cloud. The binary acquisition failed with a status buffer error message indicating target not found. A corrective maneuver as uplinked to the spacecraft, but the resulting spectrum had low counts. All the spectral observations had very low counts. The OSS could not tell if the target was in the aperture. 1.14 Completed GHRS/2 3859 (Nighttime Measurements of Nitric Oxide - Cont. of 2544) The High Resolution Spectrograph was used to make two spectral observations of bright O Class star HD116657. It is hoped to learn more about the diffusion in the Earth's upper atmosphere by studying UV absorption bands produced by nitric oxide along the line to the target star. The first spectral observations showed good counts, and the second did not. There were a couple of short slews that were executed between the observations, which pushed the target out of the aperture. 1.15 Attempted FOC/96 4804 (UV Images of Nearby Galaxies) The spacecraft entered internal hold Safemode before this observation was initiated. The observations did not take place. (See Note 2.3) 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 9 Scheduled Reacquisitions : 20 Successful : 20 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 26 Successful : 26 2.3 Significant HSTARs: HST entered Inertial Hold Safemode at 165/001726Z after failing the Earth Bright Object Protection safemode test. The system response to the entry was normal. Payload soft safing also occurred as expected. The FOC, FOS, and HSP were commanded to their safe modes while the GHRS is in hold mode and the WFPC is in ready mode. The event occurred during testing with the internal source. The safemode test failed because the FGS high voltage was prematurely on prior to exiting earth occultation. The problem appears to be a previously undiscovered error in the PASS Mission Scheduling software. Plans are to recover from safemode this afternoon. A "Health and Safety" SMS will began at 166/0800Z. Interception of the science SMS is scheduled for 167/0000Z. Internal simulations of types 11 and 3 scheduled for 6/15 and 6/16 respectively, have been canceled due to the safemode event. 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete S/C 4816 (FOC/96 Re-activation (Camera Section Only)) This item will not be executed due to the HST safemede event. 3.2 Complete HSP/POL 4888 (Remnant Stars in supernova Remnants _Aug. - Cycle 3 High) This item will not be executed due to the HST safemede event. 3.3 Complete WFC 4929 (WF/PC Cycle 2 : Improved Dark Calibration) This item will not be executed due to the HST safemede event. 3.4 Complete Two Sets of PC 3763 (Synoptic Monitoring of Seasonal Phenomena on Mars) This item will not be executed due to the HST safemede event. 3.5 Complete PC 3603 (Snapshot of Protoplanetary Nebulae) This item will not be executed due to the HST safemede event. 3.6 Complete HSP/UV2 3798 (A Search for Radial Pulsations in White Dwarfs) This item will not be executed due to the HST safemede event. 3.7 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.8 Complete FOS/RD 4504 (The Inner Regions of Quasars Cycle 3) This item will not be executed due to the HST safemede event. 3.9 Complete FOC/96 4969 (Search of QSOs Suitable for Subsequent Observations O He II) This item will not be executed due to the HST safemede event. 3.10 Complete FOC/96 4843 (Bulge Stellar Population in S0 Class Galaxies - Cycle 3 High) This item will not be executed due to the HST safemede event. 3.11 Complete HSP/UV2 4946 (Variability of High Luminosity Stars - Cycle 2 High) This item will not be executed due to the HST safemede event. 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: (See Note 2.3) 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: (See Note 2.3) /DMH -END-